OpenStudy (anonymous):

What is the product of (7.28 x 10^-2) x (9.1 x 10^-2) ? (Please express your answer in scientfic notation.) A: 6.6248 x 10^-3 B: 66.248 x 10^-4 C: 6.6248 x 10^4 D: 66.248 x 10^-3 My Answer: D

OpenStudy (anonymous):

Answer is A and B : 6.6248*10^-3 and 66.248 x 10^-4 Since 7.28*10^-2*9.1*10^-3= 7.28*9.1*10^-4 = 66.248x 10^-4=6.6248*10^-3

OpenStudy (anonymous):

ok thnx @shivamsinha